OpenStudy (anonymous):

Which is the best estimate of x based on rounding the constants and coefficients in the equation to the nearest integer? 6.15(x – 8.86) = 21.83 A. 12.8 B. 12.7 C. 12.6 D. 11.7

OpenStudy (anonymous):

@ranga @Xmoses1 @Hero @linn99123 any help if your not busy

OpenStudy (ranga):

First let us round the constants and coefficients in the equation to the nearest integer: 6.15(x – 8.86) = 21.83 after rounding becomes: 6(x - 9) = 22 divide both sides by 6 (x-9) = 22/6 = 11/3 add 9 to both sides: x = 9 + 11/3 = (27 + 11)/3 = 38/3 = 12 & 2/3 = 12.6666 rounded to 12.7

OpenStudy (anonymous):

thank you so much!

OpenStudy (ranga):

you are welcome.